Q\Determine the following ODEs are exact or not? Why?
(x/2)(cosy)y´=-1
1
Expert's answer
2020-10-27T18:56:14-0400

Given "\frac{x}{2}\cos(y)y'=-1" which can be expressed as

"\frac{x}{2}\cos(y)\frac{dy}{dx}=-1"

Equivalently we have

"\frac{x}{2}\cos(y)dy + dx=0"

The ODE is not exact since "\frac{\partial(\frac{x}{2}\cos (y))}{\partial x}=\frac{1}{2}\cos(y) \ne \frac{\partial(1)}{\partial y}=0"